Let G=fg:R--li:g(x)=ax+b,a,b EQ,a# 01. Check whether or not G is a group with respect to the composition of mappings. For f(x) = 2x + 3, find all g bilong toG such that fog=gof


1
Expert's answer
2021-07-19T15:41:06-0400

Let "f,g,h\in G" where


"f(x)=a_1x+b_1,g(x)=a_2x+b_2, \\h(x)=a_3x+b_3,a_1,b_1,a_2,b_2,a_3,b_3\in \mathbb{R}"

Consider


"f\circ g(x)=a_1(a_2x+b_2)+b_1=a_1a_2x+a_1b_2+b_1"

Clearly, "a_1a_2,a_1b_2+b_1\in \mathbb{R}" since the sum of product of real numbers is also a real number. Hence "f\circ g\in G".

Next,


"g\circ h(x)=a_2(a_3x+b_3)+b_2=a_2a_3x+a_2b_3+b_2"

So consider

"f\circ(g\circ h)(x)=a_1(a_2a_3x+a_2b_3+b_2)+b_1 \\=a_1a_2a_3x+a_1a_2b_3+a_1b_2+b_1 \\=a_1a_2(a_3x+b_3)+a_1b_2+b_1 \\=(f\circ g)\circ h(x)"

Next, let "e\in G" such that "e(x)=x". Then we have

"g\circ e(x)=a_2(e(x))+b=a_2x+b_2=g(x)=a_2x+b_2=e(g(x)) \\=e\circ g(x)"

Finally, let "k(x)=\dfrac{x}{a_2}-\dfrac{b_2}{a_2}" where "a\neq 0". Clearly, "k\in G". Now consider


"g\circ k(x)=a_2\left(\dfrac{x}{a_2}-\dfrac{b_2}{a_2}\right)+b_2=x-b_2+b_2=x \\=e(x)"

Similarly,


"k\circ g(x)=\dfrac{a_2x+b_2}{a_2}-\dfrac{b_2}{a_2}=x=e(x)"

Since "G" satisfies all the conditions of a group, we therefore conclude that "G" is a group.



Let "g(x)=ax+b" where "a,b\in \mathbb{R}". Suppose "f\circ g=g\circ f". Then we have


"2(g(x))+3=g(2x+3) \\
\Rightarrow 2(ax+b)+3=a(2x+3)+b \\
\Rightarrow 2ax+2b+3=2ax+3a+ab"

By comparing the terms of both sides of the equation above, we have


"2b+3=3a+ab \Rightarrow a(b+3)=2b+3 \\ \Rightarrow a=\dfrac{2b+3}{b+3}"

where "b\neq -3".

Hence we have the set

"\{f \in G | f(x)=ax+b \text{ where }a=\dfrac{2b+3}{b+3},b \neq -3\}"